OpenStudy (anonymous):

sketch the region bounded by the line y=x and the parabola y=6-x^2

OpenStudy (anonymous):

and find its area

OpenStudy (amistre64):



OpenStudy (amistre64):

the height of each "rect. area" is (6-x^2)-(x)

OpenStudy (anonymous):

thanks, I know y=x but i dnt usually know how to find the parabola. can you tell me how you go about that. and also the froula for area is (integreal sign)ydx

OpenStudy (amistre64):

so all we have to do is add up all the heighths*width rect.s from endpoint to endpoint

OpenStudy (amistre64):

y=x^2 is your usualy granddaddy parabola; all the parabolas are an alteration of him

OpenStudy (amistre64):

when we want to flip x^2 over, we simply place a "-" in front y=-x^2 to move it up by 6 ...well, we add 6 y = -x^2+6 ; and we can write it as: y = 6-x^2 for kicks as well

OpenStudy (anonymous):

thank much, it make sperfect sense

OpenStudy (amistre64):